OpenStudy (anonymous):

Find the distance between the points with given coordinates. (w - 2, w) and (w, 4w) a) √(4 + 3w^2) b) √(4 + 9w^2) c) √(-4 + 9w^2) d) √(-4 + 3w^2)

OpenStudy (nali):

distance formula d = \[\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\]

OpenStudy (nali):

squreroot (w - w-2)^2 + (4w-w)^2 w-w = 0 so it would be (-2)^2 = 4 and 4w-w = 3w then (3w)^2 = 9w^2 so the answer is squreroot 4+ 9w^2 ____> b)